# Generative Ontology: When Structured Knowledge Learns to Create

### v1

@@ Abstract
Generative Ontology encodes domain knowledge as executable Pydantic schemas that constrain LLM generation via DSPy signatures. A multi-agent pipeline assigns specialized roles to different ontology domains: a Mechanics Architect designs game systems, a Theme Weaver integrates narrative, a Balance Critic identifies exploits. Each agent carrying a professional “anxiety” that prevents shallow, agreeable outputs. Retrieval-augmented generation grounds novel designs in precedents from existing exemplars, while iterative validation ensures coherence between mechanisms and components.
We demonstrate the framework through GameGrammar, a system for generating complete tabletop game designs. Given a thematic prompt (“bioluminescent fungi competing in a cave ecosystem”), the pipeline produces structurally complete, playable game specifications with mechanisms, components, victory conditions, and setup instructions. These outputs satisfy ontological constraints while remaining genuinely creative.
The pattern generalizes beyond games. Any domain with expert vocabulary, validity constraints, and accumulated exemplars (music composition, software architecture, culinary arts) is a candidate for Generative Ontology. We argue that constraints do not limit creativity but enable it: just as grammar makes poetry possible, ontology makes structured generation possible.
@@ 1.2  Complementary Limitations
1.3  Related Work
@@
Mehenni and Zouaq [12] demonstrate ontology-constrained generation for clinical summaries, the closest methodological precedent to our work. However, their goal is hallucination reduction in summarization, whereas we target structured creativity. In game generation, Gallotta et al. [9] survey LLM applications in games broadly, while Becker et al. [4] present Boardwalk, a framework for generating board game code from existing rules, generating implementations, not designs. Maleki and Zhao [11] survey procedural content generation methods, noting emerging LLM integration but without ontological constraints. Cooper [7] provides foundational work on constrained decoding that underpins our structured output approach.
Our contribution fills a specific gap: using ontology as a generative grammar for structured creative output, operationalized through executable schemas, multi-agent pipelines, and retrieval augmentation.
1.4  Paper Overview
We introduce Generative Ontology and develop it through the domain of tabletop game design. Section 2 defines the framework and its philosophical grounding. Section 3 transforms ontology classes into executable Pydantic schemas. Section 4 operationalizes generation with DSPy signatures. Section 5 introduces multi-agent specialization with professional anxieties. Section 6 adds retrieval-augmented grounding. Section 7 develops the ontology as a validation contract. Section 8 presents a complete case study. Section 9 argues the pattern generalizes beyond games. Section 10 concludes with future directions.
@@ 2.2  Philosophical Grounding
For readers familiar with our earlier exploration of process philosophy for AI agent design [6], Generative Ontology has a natural grounding in Alfred North Whitehead’s metaphysics [16]. Whitehead distinguished between eternal objects (pure forms, patterns, and structures that exist as potentials) and actual occasions (the concrete events where eternal objects find expression). Our game ontology is a collection of eternal objects: the abstract patterns of worker placement, deck building, area control. These patterns have no reality until they are instantiated in actual games: in Catan’s modular hex board, in Dune: Imperium’s faction tracks. Generative Ontology is the process of moving from eternal objects to actual occasions: the ontology provides the forms; generation produces the instances. Creativity, in Whitehead’s view, is precisely this novel synthesis of eternal objects into actual occasions that have never existed before [3]. The forms constrain but do not determine. Each generated game is a genuine novelty.
@@ 3.3  Why Pydantic?
The original ontology was a tool for analysis: we could take Catan and decompose it into ontology terms. The schema flips this relationship. Now the ontology becomes a tool for synthesis: we provide a theme, and the constrained generation produces a complete game that satisfies every ontological requirement.
@@ 4.1  Signatures as Ontological Contracts
Listing 2: DSPy GenerativeOntologySignature. The typed output field binds generation to the ontological schema.
@@ 4.2  Execution Flow
DSPy’s ChainOfThought module extends this by instructing the model to reason through the design before committing to structured output. The reasoning captures why the system chose worker placement over deck building, how the theme informed the end condition, and what creates tension in the player dynamics. This reasoning is not merely debugging output; it documents the design rationale.
Encapsulating logic in a DSPy Module enables composition of multiple generation steps, post-generation validation, and retry logic. The module adds a layer of semantic validation beyond what Pydantic provides: the schema ensures structural correctness (all fields present, correct types), while the module ensures coherence: that mechanisms match components and that game type aligns with player dynamics. This layered architecture (schema for structure, module for semantics) is central to the reliability of Generative Ontology.
@@ 5  Multi-Agent Ontology Pipeline
A single LLM call, even with our ontology schema, carries the entire burden of game design: it must simultaneously consider mechanisms, theme, components, balance, and player experience. Human design teams do not work this way. A designer sketches mechanics while an artist develops visual language while a playtester identifies broken interactions. Generative Ontology enables the same division of labor by decomposing the ontology into domains and assigning specialized agents to each.
@@ 5.1  Agent Roster and Professional Anxieties
Table 4 lists the five agents. The “Anxiety” column is crucial: each agent is not merely assigned a domain but is given a concern, a professional worry that shapes its generation and critique. The Mechanics Architect does not just produce mechanisms; it worries about whether those mechanisms create meaningful choices. This anxiety-driven design prevents the “yes-man” tendency of LLMs to produce plausible but shallow output.
@@
Each agent receives a DSPy signature tailored to its domain, with the docstring encoding its persona and anxiety. The Mechanics Architect’s signature begins: “You are a veteran game mechanics designer who has studied hundreds of tabletop games. You obsess over player agency…” The Balance Critic’s opens: “You are a competitive player and game breaker. Your job is to find what’s broken…” These personas are not decorative; they shape the distribution of generated outputs toward domain-appropriate rigor.
@@ 5.2  Pipeline Architecture
    Sequential Generation. The Mechanics Architect produces core systems given the theme and constraints. The Theme Weaver receives these mechanics and integrates narrative. The Component Designer receives both and specifies physical elements. Each agent builds on the previous, with information flowing through typed Pydantic schemas.
@@
    Critical Evaluation. The Balance Critic examines the assembled design for exploits, dominant strategies, and interaction gaps. Issues are tagged with severity levels (minor, moderate, severe).
@@
    Experience Assessment. The Fun Factor Judge evaluates engagement hooks, tension moments, satisfaction sources, and replayability, producing a 1–10 rating and improvement suggestions.
5.3  Why Multi-Agent Outperforms Single-Pass
The multi-agent approach offers five advantages: (1) focused expertise, where each agent optimizes for its domain without distraction; (2) built-in tension, where agents have different anxieties, mirroring real design team dynamics; (3) explicit handoffs, where information flows through defined schema interfaces; (4) iterative refinement, where critics identify issues that refiners address; and (5) observability, where when something fails, we know which agent to adjust.
@@ 6  Retrieval-Augmented Generation
Our Generative Ontology can create novel game designs, but creativity without context risks reinventing solutions that already exist, or repeating mistakes that experienced designers have learned to avoid. The most interesting creative work happens not in a vacuum but in conversation with what has come before.
Figure 4: Knowledge System architecture. A unified Retriever Service dispatches queries to ChromaDB (semantic vector search) or SQLite (structured fallback), drawing on a BGG corpus of 2,000 games with 384-dimensional embeddings, mechanism types, and filterable metadata.
@@ 6.1  Two-Phase Retrieval
As a concrete example, for the prompt “trading empires in the ancient Mediterranean” with target mechanisms [TRADING, AREA_CONTROL, RESOURCE_MANAGEMENT], the system retrieves Concordia (card-driven trading networks across the Roman Empire), Cyclades (Greek island control with auction-based resource competition), and Tigris & Euphrates (civilization building with multi-dimensional area control scoring). The generated design synthesizes insights from Concordia’s card mechanism and Tigris & Euphrates’ balanced scoring into something novel that explicitly acknowledges its lineage in a design rationale field.
@@ 6.2  Grounded Creativity
RAG transforms Generative Ontology from invention to informed invention. The system knows what mechanisms have been successfully combined, how similar themes have been approached, and what design patterns work for given player counts and complexity levels. When retrieval returns few results for a mechanism–theme combination, we have discovered an unexplored design space. The absence of precedent is itself information. The ontology provides the grammar; the retrieved examples provide the vocabulary. Together, they enable generation that speaks the language of game design fluently.
@@ 7  Ontology as Validation Contract
Generation is not enough. LLMs are notoriously agreeable: they produce output that looks valid without ensuring it is valid. The ontology serves not only as a generation schema but as a validation contract: a set of constraints that generated designs must satisfy to count as coherent.
7.1  The Yes-Man Problem
An LLM asked to generate a deck-building game may mention deck_building as a mechanism yet include no cards in the components, producing structurally incoherent output that sounds like a deck-building game without being one. Pydantic validation catches type errors and missing fields, but semantic coherence (the relationship between mechanisms and components) requires deeper validation.
@@ 7.3  Iterative Refinement
DSPy’s Assert mechanism enables inline validation: when dspy.Assert(is_valid, error_message) fails, DSPy automatically retries generation with the error message included as feedback. For complex validation requiring multiple attempts, we implement an explicit refinement loop: generate, validate, refine with specific issue descriptions, and repeat up to a configurable maximum. The validator returns both a boolean and a list of specific issues, enabling targeted correction rather than wholesale regeneration.
@@
This architecture treats the ontology as a contract between the generation system and downstream consumers. Downstream systems (whether human designers reviewing output, automated balancing tools, or game engines instantiating the design) can rely on structural guarantees: a deck-building game will have cards; a cooperative game will not have direct conflict; the end condition will be specified.
@@ 8  Case Study: Mycelium: The Deep
Theory is persuasive; demonstration is convincing. We trace a complete generation through the Generative Ontology pipeline, from initial theme to playable game design.
8.1  Input
We provide a theme chosen to be unusual enough that retrieval finds no direct precedents, forcing the system to synthesize rather than imitate:
@@ 8.1  Input
8.2  Phase 1: Mechanics Architect
The Mechanics Architect analyzes the theme and selects three core mechanisms: RESOURCE_MANAGEMENT (spore allocation across growth opportunities), AREA_CONTROL (territory expansion through the cave network), and ENGINE_BUILDING (specialized fungal structures providing ongoing benefits). The turn structure uses simultaneous planning (players place spore tokens secretly) followed by sequential resolution. The agency justification notes that players must balance short-term growth against long-term network health, with engine choices creating divergent strategies even in cooperative play.
8.3  Phase 2: Theme Weaver
The Theme Weaver integrates narrative, titling the game Mycelium: The Deep. The setting is an ancient limestone cave system where bioluminescent fungi are the only light source, slowly flooding as underground rivers shift. Thematic integration maps each mechanism to its fictional counterpart: spore placement is reproductive strategy; area control is securing nutrient pathways; engine building is developing specialized structures (fruiting bodies for points, rhizomorphs for fast movement, symbiotic bacteria for defense). The antagonist is the rising water table, creating cooperative tension without player-versus-player conflict.
8.4  Phase 3: Component Designer
The Component Designer specifies 30 modular hex cave tiles arranged vertically (deeper at the bottom), three card types (Event, Structure, Objective), example cards (“Flash Flood: Water rises 2 levels this round”; “Fruiting Body: Costs 3 growth, produces 1 VP per round”), and tokens including fungal growth cubes (15 per player in 4 colors), spore tokens (3 per player, regenerating each round), and flood markers (20 blue cubes advancing upward).
Figure 5: Case study: the design conversation. Each agent builds on the previous, with critics providing feedback that drives refinement before final assessment.
8.5  Phase 4: Balance Critic
The Balance Critic identifies three issues. First, a runaway leader problem in cooperative play: if one player’s strain grows faster, they dominate decisions (severity: moderate; recommendation: add a network synergy rule so connected strains share benefits). Second, arbitrary flood timing: random event cards determining flood speed leaves players with no agency over the core tension (severity: moderate; recommendation: make flood a predictable escalation of one level per round, with events as modifiers). Third, engine building as a potential trap: structures costing growth may cause builders to fall behind (severity: minor; recommendation: structures should provide an immediate small benefit plus an ongoing ability).
8.6  Phase 5: Refinement
The refinement agent addresses both moderate-severity issues: it adds a Network Synergy rule (adjacent strains share engine abilities), changes flood timing to a predictable one-level-per-round baseline with event modifiers, and rebalances structures to provide an immediate 1 VP plus their ongoing ability.
8.7  Phase 6: Fun Factor Judge
The Fun Factor Judge rates the refined design 7/10. Engagement hooks include persistent tension (“Will we reach the surface in time?”), simultaneous placement creating reveal anticipation, and structure building as achievement. Tension moments include the visible flood countdown, event card reveals, and decisions about saving endangered colonies versus abandoning them. Replayability stems from modular board setup, different strain combinations ($4!=24$ team configurations), and event deck variability.
8.8  Output Summary
Table 5 presents key fields of the generated design.
@@ 8.8  Output Summary
Players 2–4, role-differentiated (Spreader, Rooted, Lumens, Toxic)
@@
8.9  Analysis
Without the ontology schema, an LLM generating “a cooperative game about bioluminescent fungi” would likely produce vague victory conditions (“escape the cave somehow”), inconsistent mechanisms (mentioning deck-building without specifying cards), missing components, unbalanced elements, and disconnected theme. Generative Ontology enforced complete goal specification via minimum-length constraints, coherent mechanism–component alignment via the validation function, thematic integration via the Theme Weaver’s explicit mapping, playability basics via required fields for turn structure and end condition, and balance review via the critic agent’s adversarial posture. The output is not merely plausible; it is playable. A designer could take this output, build a prototype, and begin playtesting immediately.

### v2

@@ Abstract
Generative Ontology encodes domain knowledge as executable Pydantic schemas that constrain LLM generation via DSPy signatures. A multi-agent pipeline assigns specialized roles to different ontology domains: a Mechanics Architect designs game systems, a Theme Weaver integrates narrative, a Balance Critic identifies exploits, each carrying a professional “anxiety” that prevents shallow, agreeable outputs. Retrieval-augmented generation grounds novel designs in precedents from existing exemplars, while iterative validation ensures coherence between mechanisms and components.
We demonstrate the framework through GameGrammar, a system for generating complete tabletop game designs, and present three empirical studies. An ablation study (120 designs, 4 conditions) shows that multi-agent specialization produces the largest creative quality gains (fun $d=1.12$, strategic depth $d=1.59$; both $p<.001$), while schema validation eliminates structural errors ($d=4.78$). A benchmark comparison against 20 published board games reveals structural parity but a bounded creative gap (fun $d=1.86$): generated designs score 7–8 while published games score 8–9, with tension/drama and social interaction already at parity. A test-retest reliability study (ICC analysis, 50 evaluations) validates the LLM-based evaluator, with 7 of 9 metrics achieving Good-to-Excellent reliability (ICC 0.836–0.989).
The pattern generalizes beyond games. Any domain with expert vocabulary, validity constraints, and accumulated exemplars is a candidate for Generative Ontology. Constraints do not limit creativity but enable it: just as grammar makes poetry possible, ontology makes structured generation possible.
@@ 1.2  Complementary Limitations
1.3  Paper Overview
We introduce Generative Ontology and develop it through the domain of tabletop game design. Section 2 reviews related work. Section 3 defines the framework and its philosophical grounding. Sections 48 develop the technical architecture: executable schemas, DSPy operationalization, multi-agent pipelines, retrieval augmentation, and validation contracts. Section 9 presents three empirical studies evaluating the framework. Section 10 illustrates the pipeline with a complete case study. Section 11 argues the pattern generalizes beyond games. Section 13 concludes with future directions.
2  Related Work
2.1  Ontologies and LLMs
@@
Mehenni and Zouaq [13] demonstrate ontology-constrained generation for clinical summaries, the closest methodological precedent to our work, though their goal is hallucination reduction in summarization rather than structured creativity. Cooper [7] provides foundational work on constrained decoding that underpins our structured output approach.
2.2  LLMs in Game Design
Gallotta et al. [9] survey LLM applications in games broadly, while Becker et al. [4] present Boardwalk, a framework for generating board game code from existing rules, generating implementations rather than designs. Maleki and Zhao [12] survey procedural content generation methods, noting emerging LLM integration but without ontological constraints.
2.3  LLM-as-Judge Evaluation
Recent work has established LLMs as evaluators of creative output [20]. Concerns about evaluator reliability motivate our test-retest analysis: Shankar et al. [16] argue that “who validates the validators” is a critical question for LLM-based evaluation pipelines. We address this by reporting intraclass correlation coefficients (ICC) for all evaluation metrics [11, 17], demonstrating that our evaluator is a reliable instrument rather than assuming so.
2.4  Gap
Our contribution fills a specific gap: using ontology as a generative grammar for structured creative output, operationalized through executable schemas, multi-agent pipelines, and retrieval augmentation, with rigorous empirical evaluation including ablation, benchmarking, and evaluator validation.
@@ 3.2  Philosophical Grounding
For readers familiar with our earlier exploration of process philosophy for AI agent design [6], Generative Ontology has a natural grounding in Alfred North Whitehead’s metaphysics [19]. Whitehead distinguished between eternal objects (pure forms, patterns, and structures that exist as potentials) and actual occasions (the concrete events where eternal objects find expression). Our game ontology is a collection of eternal objects: the abstract patterns of worker placement, deck building, area control. These patterns have no reality until they are instantiated in actual games. Generative Ontology is the process of moving from eternal objects to actual occasions: the ontology provides the forms; generation produces the instances. Creativity, in Whitehead’s view, is precisely this novel synthesis of eternal objects into actual occasions that have never existed before [3].
@@ 5.1  Signatures as Ontological Contracts
Listing 2: DSPy generation signature. The typed output field binds generation to the ontological schema.
@@ 5.2  Execution Flow
DSPy’s ChainOfThought module extends this by instructing the model to reason through the design before committing to structured output. Encapsulating logic in a DSPy Module enables composition of multiple generation steps, post-generation validation, and retry logic. The module adds a layer of semantic validation beyond what Pydantic provides: the schema ensures structural correctness (all fields present, correct types), while the module ensures coherence: that mechanisms match components and that game type aligns with player dynamics.
@@ 6  Multi-Agent Ontology Pipeline
A single LLM call, even with our ontology schema, carries the entire burden of game design: it must simultaneously consider mechanisms, theme, components, balance, and player experience. Human design teams do not work this way. Generative Ontology enables the same division of labor by decomposing the ontology into domains and assigning specialized agents to each.
@@ 6.1  Agent Roster and Professional Anxieties
Table 4 lists the five agents. The “Anxiety” column is crucial: each agent is not merely assigned a domain but is given a concern, a professional worry that shapes its generation and critique. This anxiety-driven design prevents the “yes-man” tendency of LLMs to produce plausible but shallow output.
@@ 6.2  Pipeline Architecture
    Sequential Generation. The Mechanics Architect produces core systems given the theme and constraints. The Theme Weaver receives these mechanics and integrates narrative. The Component Designer receives both and specifies physical elements.
@@
    Critical Evaluation. The Balance Critic examines the assembled design for exploits, dominant strategies, and interaction gaps.
@@
    Experience Assessment. The Fun Factor Judge evaluates engagement hooks, tension moments, satisfaction sources, and replayability.
@@ 7  Retrieval-Augmented Generation
Our Generative Ontology can create novel game designs, but creativity without context risks reinventing solutions that already exist. The most interesting creative work happens in conversation with what has come before.
Figure 4: Knowledge System architecture. A unified Retriever Service dispatches queries to ChromaDB (semantic vector search) or SQLite (structured fallback), drawing on a BGG corpus of 1,767 games with 384-dimensional embeddings, mechanism types, and filterable metadata.
@@ 7.2  Grounded Creativity
RAG transforms Generative Ontology from invention to informed invention. The system knows what mechanisms have been successfully combined, how similar themes have been approached, and what design patterns work for given player counts and complexity levels. The ontology provides the grammar; the retrieved examples provide the vocabulary. Together, they enable generation that speaks the language of game design fluently.
@@ 8  Ontology as Validation Contract
Generation is not enough. The ontology serves not only as a generation schema but as a validation contract: a set of constraints that generated designs must satisfy to count as coherent.
@@ 8.2  Iterative Refinement
DSPy’s Assert mechanism enables inline validation: when dspy.Assert(is_valid, error_message) fails, DSPy automatically retries generation with the error message included as feedback. This architecture treats the ontology as a contract between the generation system and downstream consumers. Downstream systems (whether human designers, balancing tools, or game engines) can rely on structural guarantees: a deck-building game will have cards; a cooperative game will not have direct conflict; the end condition will be specified.
9  Empirical Evaluation
The preceding sections describe the Generative Ontology framework. We now evaluate it through three complementary studies: an ablation study isolating the contribution of each architectural component, a benchmark comparison against published board games, and a test-retest reliability analysis validating our evaluation instrument.
9.1  Experimental Setup
Generation Model. All designs were generated using Claude Sonnet 4 (model ID claude-sonnet-4-20250514, temperature 0.7) via DSPy.
Evaluation Pipeline. Each generated design is evaluated through two complementary lenses: (1) structural metrics, computed deterministically by checking the ontology for completeness (are all required fields present?), alignment (do mechanisms match components?), and consistency (are there contradictions?); and (2) creative metrics, scored by an LLM evaluator (“Design Coach”) on seven dimensions: strategic depth, tension/drama, player agency, replayability, social interaction, elegance, and thematic cohesion, each on a 1–10 scale. The Design Coach also provides an overall fun rating (1–10) and engagement variance (standard deviation across dimension scores). The evaluator uses the same model at temperature 0 for near-deterministic scoring.
Theme Prompts. Ten standardized theme prompts span a range of complexity levels and game types (see Appendix A). Each prompt specifies a theme, player count, complexity level, and target play time.
Experimental Conditions. Four conditions isolate the contribution of each pipeline component (Table 5).
Table 5: Experimental conditions for the ablation study.
Condition Configuration
C1 (Baseline) Raw LLM generation, no schema or ontology
C2 (Schema) Pydantic schema validation only
C3 (Ontology) Schema + ontology constraints, single agent
C4 (Pipeline) Full pipeline: schema + ontology + RAG + multi-agent + validation
9.2  Study 1: Ablation
Design. $4\text{ conditions}\times 10\text{ prompts}\times 3\text{ replications}=120$ designs, all evaluated through the full pipeline.
[…]
Table 6: Structural metrics by condition.
Condition Errors Compl.
C1 (Baseline) 5.03 100%
C2 (Schema) 0.10 100%
C3 (Ontology) 0.00 83%
C4 (Pipeline) 0.00 100%
Figure 5: Structural metrics by condition. Left: mean consistency errors drop from 5.03 (C1) to zero (C3, C4). Right: completeness is 100% for all conditions except C3 (83%), where single-agent generation misses detail fields.
[…]
Figure 6: Creative dimension score distributions by condition. C4 (Pipeline) shows the highest and tightest distributions on most dimensions. Replayability and social interaction do not discriminate between conditions.
[…]
Figure 7: Cost analysis. Left: generation tokens per design. Right: tokens per fun point. C4 costs ${\sim}8.5\times$ more but achieves the highest fun rating.
Summary. Structural validity is solved by schema validation. Creative quality arises from the interaction between constraint enforcement and architectural specialization: constraints alone suppress richness; specialization restores and enhances it.
9.3  Study 2: Benchmark Comparison
Design. We compare 30 C4-generated designs from Study 1 against 20 published board games evaluated through the same pipeline. The reference set spans three complexity tiers: light (Azul, Splendor, Love Letter, Kingdomino, Codenames), medium (Wingspan, Everdell, 7 Wonders, Concordia, Spirit Island), and heavy (Brass: Birmingham, Gaia Project, Twilight Imperium 4E, Through the Ages, Gloomhaven), plus five pre-existing ontologies (Catan, Root, Dune Imperium, Terraforming Mars, Ticket to Ride). See Appendix B for the full list. Each real game was reverse-engineered into a complete GameGrammar ontology and evaluated through the identical Design Coach pipeline.
Structural Parity. Both groups achieve 100% completeness. Generated designs have fewer consistency errors (1.27 vs. 2.80, $d=0.76$, $p<.01$). The higher error count for real games reflects richer component vocabularies (specialized tokens, standees, custom boards) that extend beyond the checker’s standard categories.
Creative Gap. Table 7 reports creative dimension scores. Published games score significantly higher on 5 of 7 dimensions, with the largest effects on elegance ($d=1.72$) and fun ($d=1.86$). Crucially, tension/drama ($d=0.35$, ns) and social interaction ($d=0.19$, ns) show no significant differences, suggesting the pipeline already handles these dimensions comparably.
Table 7: Creative dimensions: real vs. generated designs.
Dimension Real Gen. $d$ Sig.
Strategic Depth 8.85 8.13 0.94 **
Tension & Drama 8.50 8.20 0.35 ns
Player Agency 9.00 8.23 1.21 ***
Replayability 9.05 7.63 1.59 ***
Social Interaction 7.20 6.93 0.19 ns
Elegance 9.25 7.97 1.72 ***
Fun Rating 8.90 8.07 1.86 ***
$d$ = Cohen’s $d$; ns = not significant; ** $p<.01$; *** $p<.001$.
Figure 8: Violin plots showing score distributions for all creative dimensions. Real games (red) cluster in the 8–9 range while generated designs (green) spread across 7–9. Tension/drama and social interaction show the most overlap.
Mechanism Similarity. Generated designs use a similar number of mechanisms (6.7 vs. 6.1, d=0.60d=-0.60) but fewer component types (7.8 vs. 11.6, $d=1.12$, $p<.01$). The pipeline generates designs within a mid-complexity band but does not yet produce the extremes: neither ultra-light party games nor epic-scale strategy games (Figure 9).
Figure 9: Mechanism count distributions. Both groups center around 6–7 mechanisms; generated designs show slightly more spread.
Summary. Generated designs achieve structural parity with published games but exhibit a bounded creative gap of approximately one point on a 10-point scale. This gap likely reflects properties that emerge from iterative playtesting (elegance and replayability) rather than design specification alone.
9.4  Study 3: Evaluator Reliability
An empirical evaluation is only as credible as its evaluation instrument. We assess the Design Coach’s test-retest reliability using intraclass correlation coefficients (ICC).
Design. Ten designs spanning the quality range (fun ratings 6–9, all four conditions) were each evaluated five times through the Design Coach, yielding 50 evaluations. We compute ICC(2,1), two-way random effects with single rater and absolute agreement [17], for each of the nine LLM-scored metrics.
Results. Table 8 reports ICC values. Seven of nine metrics achieve Good-to-Excellent reliability (ICC $\geq 0.75$). Four metrics rate Excellent ($\geq 0.90$): tension & drama (0.989), social interaction (0.983), thematic cohesion (0.975), and elegance (0.934). Three rate Good: fun rating (0.873), strategic depth (0.846), and player agency (0.836). Two metrics, replayability (0.578) and engagement variance (0.598), show moderate reliability.
Table 8: Test-retest reliability: ICC(2,1) per metric.
Metric ICC 95% CI Level
Tension & Drama .989 [.970, 1.00] Excellent
Social Interaction .983 [.960, .990] Excellent
Thematic Cohesion .975 [.940, .990] Excellent
Elegance .934 [.850, .980] Excellent
Fun Rating .873 [.730, .960] Good
Strategic Depth .846 [.680, .950] Good
Player Agency .836 [.660, .950] Good
Engagement Var. .598 [.330, .850] Moderate
Replayability .578 [.300, .840] Moderate
Figure 10: ICC values per metric with 95% confidence intervals. Green = Excellent ($\geq.90$), blue = Good (.75–.90), orange = Moderate (.50–.75).
Cross-Study Validation. The ICC results retroactively validate prior findings. Fun rating (ICC = 0.873) confirms that the fun-based comparisons in Studies 1 and 2 rest on stable measurements. The high ICC for tension & drama (0.989) and social interaction (0.983) confirms that the non-significant differences in Study 2 reflect genuine equivalence, not measurement noise. Conversely, replayability’s moderate ICC (0.578) suggests that its non-significance in Study 1 may partly reflect evaluator inconsistency, a property arguably better assessed through playtesting than design specification analysis.
Summary. The Design Coach is a reliable evaluation instrument. Score differences between designs reflect genuine quality differences, not evaluator drift.
@@ 10  Case Study: Mycelium: The Deep
To illustrate the pipeline concretely, we trace a single generation from theme to playable design. The input is a theme chosen to be unusual enough that retrieval finds no direct precedents:
@@
The Mechanics Architect selects three core mechanisms: resource management (spore allocation), area control (cave network expansion), and engine building (specialized fungal structures). The Theme Weaver titles the game Mycelium: The Deep and maps each mechanism to its fictional counterpart: spore placement as reproductive strategy, area control as securing nutrient pathways, engine building as developing specialized structures. The Component Designer specifies 30 modular hex cave tiles, three card types, and tokens including fungal growth cubes, spore tokens, and flood markers. The Balance Critic identifies a runaway-leader problem and arbitrary flood timing; the refinement agent adds a Network Synergy rule and predictable flood escalation. The Fun Factor Judge rates the refined design 7/10. Table 9 summarizes the output.
@@
Players 2–4, role-differentiated
@@
Without the ontology schema, an LLM generating “a cooperative game about bioluminescent fungi” would likely produce vague victory conditions, inconsistent mechanism–component mappings, and disconnected theme. Generative Ontology enforced complete goal specification, coherent mechanism–component alignment, and thematic integration. The output is not merely plausible; it is playable.